OpenStudy (anonymous):

John has a total of 13 coins in his pocket. Of the 13 coins, they are all either quarters or dimes. The total value of the coins is $2.50. How many quarters does John have in his pocket?

OpenStudy (anonymous):

By easy thinking - since the last digit is 0, he needs to have even quarters. 10 quarters is too much (in total, he would have over 2.50), 8 quarters is just enough (8 quarters are $2 + 5 dimes and it's $2.50)

OpenStudy (anonymous):

so the answer is 8 qurarters right

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

q + d = 13 --> d = 13 - q .25q + .10d = 2.50 now sub 13 - q in for d in the 2nd equation .25q + .10(13 - q) = 2.50 .25q + 1.30 - .10q = 2.50 .15q = 2.50 - 1.30 .15q = 1.20 q = 1.20/.15 q = 8 now sub 8 in for q in the 1st equation q = d = 13 8 + d = 13 d = 13 - 8 d = 5 check... .25q + .10d = 2.50 .25(8) + .10(5) = 2.50 2.00 + .50 = 2.50 2.50 = 2.50 (correct) There are 8 quarters and 5 dimes